Solve the differential equation : (xy^2 - x^2) dx +(3x^2y^2 +x^2y-2x^3+y^2) dy =0.
1
Expert's answer
2020-04-10T17:14:29-0400

Lets rewrite the equation as below:


"\left(xy^2-x^2\right)dx+\left(3x^2y^2+x^2y-2x^3+y^2\right)dy=0\\[0.3cm]
p(x,y)dx+q(x,y)dy=0\\[0.3cm]
\left\{\begin{array}{l}
p(x,y)=xy^2-x^2\\[0.3cm]
q(x,y)=3x^2y^2+x^2y-2x^3+y^2
\end{array}\right."

The equation is called exact differential equation if it satisfied:



"\frac{\partial p(x,y)}{\partial y}=\frac{\partial q(x,y)}{\partial y}"

In our case,



"\left\{\begin{array}{l}
\displaystyle\frac{\partial p}{\partial y}=2xy\\[0.3cm]
\displaystyle\frac{\partial q}{\partial x}=6xy^2+2xy-6x^2
\end{array}\right.\longrightarrow\\[0.3cm]
\frac{\partial p}{\partial y}\neq\frac{\partial q}{\partial x}"

Since the relation is not equal, this differential equation is not exact.

Next step is find a function "u(x,y)" which will make equation 

"u(x,y)(p(x,y)dx+q(x,y)dy)=0"  become exact differential equation.

"u(x,y)"  is called integrating factor, and will be much easier to find if it is only "x" function or only "y" function.

"u(x,y)"  is only x function if



"u(x,y)\equiv u(x)=\frac{\displaystyle\frac{\partial p}{\partial y}-\displaystyle\frac{\partial q}{\partial x}}{q(x,y)}"

only have "x" variable, and it is only "y" function if



"u(x,y)\equiv u(y)=\frac{\displaystyle\frac{\partial q}{\partial x}-\displaystyle\frac{\partial p}{\partial y}}{p(x,y)}"



only have "y" variable.

In our case,



"\frac{\displaystyle\frac{\partial p}{\partial y}-\displaystyle\frac{\partial q}{\partial x}}{q(x,y)}=\frac{2xy-\left(6xy^2+2xy-6x^2\right)}{3x^2y^2+x^2y-2x^3+y^2}\\[0.3cm]
\frac{\displaystyle\frac{\partial p}{\partial y}-\displaystyle\frac{\partial q}{\partial x}}{q(x,y)}=\frac{6x\left(x-y^2\right)}{3x^2y^2+x^2y-2x^3+y^2}\\[0.3cm]"

contain both "x" and "y" .



"\frac{\displaystyle\frac{\partial q}{\partial x}-\displaystyle\frac{\partial p}{\partial y}}{p(x,y)}=\frac{\left(6xy^2+2xy-6x^2\right)-2xy}{xy^2-x^2}\\[0.3cm]
\frac{\displaystyle\frac{\partial q}{\partial x}-\displaystyle\frac{\partial p}{\partial y}}{p(x,y)}=\frac{6x\left(y^2-x\right)}{x(y^2-x)}=6\\[0.3cm]"

contain only "y" variable.

So "u(x,y)"  is only "y" function. In this case:



"u(y)=e^{\int 6dy}=e^{6y}"

Back to our initial equation:



"u(y)(p(x,y)dx+q(x,y)dy)=0\\[0.3cm]
e^{6y}\left(xy^2-x^2\right)dx+e^{6y}\left(3x^2y^2+x^2y-2x^3+y^2\right)dy=0"

Now find a function "f(x,y)"  that satisfied:



"\left\{\begin{array}{l}
\displaystyle\frac{\partial f}{\partial x}=e^{6y}\left(xy^2-x^2\right)\\[0.3cm]
\displaystyle\frac{\partial f}{\partial y}=e^{6y}\left(3x^2y^2+x^2y-2x^3+y^2\right)
\end{array}\right.\\[0.3cm]
\frac{\partial f}{\partial x}=e^{6y}\left(xy^2-x^2\right)\longrightarrow\\[0.3cm]
f(x,y)=\int\left(e^{6y}\left(xy^2-x^2\right)\right)dx+g(y)\longrightarrow\\[0.3cm]
\boxed{f(x,y)=e^{6y}\left(\frac{x^2y^2}{2}-\frac{x^3}{3}\right)+g(y)}\\[0.3cm]
\frac{\partial f}{\partial y}=\frac{\partial}{\partial y}\left(e^{6y}\left(\frac{x^2y^2}{2}-\frac{x^3}{3}\right)+g(y)\right)=\\[0.3cm]
=6\cdot e^{6y}\left(\frac{x^2y^2}{2}-\frac{x^3}{3}\right)+e^{6y}\left(x^2y\right)+\frac{dg}{dy}=\\[0.3cm]
=e^{6y}\left(3x^2y^2-2x^3+x^2y\right)+\frac{dg}{dy}"

Then,


"e^{6y}\left(3x^2y^2-2x^3+x^2y\right)+\frac{dg}{dy}=e^{6y}\left(3x^2y^2+x^2y-2x^3+y^2\right)\\[0.3cm]
\frac{dg}{dy}=e^{6y}\cdot y^2\longrightarrow g(y)=\int\left(e^{6y}y^2\right)dy\\[0.3cm]
\left(\text{integration by parts}\right) : \\[0.3cm]
g(y)=\int\left(e^{6y}y^2\right)dy=\frac{y^2e^{6y}}{6}-\int\frac{2e^{6y}}{6}dy=\\[0.3cm]
=\frac{y^2e^{6y}}{6}-\frac{1}{3}\cdot\left(\frac{ye^{6y}}{6}-\int\frac{e^{6y}}{6}\right)=\\[0.3cm]
=\frac{y^2e^{6y}}{6}-\frac{ye^{6y}}{18}-\frac{e^{6y}}{108}\equiv\frac{e^{6y}\left(18y^2-6y+1\right)}{108}"

Conclusion,



"\boxed{f(x,y)=f(x,y)=e^{6y}\left(\frac{x^2y^2}{2}-\frac{x^3}{3}\right)+\frac{e^{6y}\left(18y^2-6y+1\right)}{108}}"

And finally our equation can be rewriten as:



"\frac{\partial f(x,y)}{\partial x}\cdot dx+\frac{\partial f(x,y)}{\partial y}\cdot dy=df(x,y)=0\\[0.3cm]
f(x,y)=C\\[0.3cm]
\boxed{e^{6y}\left(\frac{x^2y^2}{2}-\frac{x^3}{3}\right)+\frac{e^{6y}\left(18y^2-6y+1\right)}{108}=C}"

ANSWER

General solution



"e^{6y}\left(\frac{x^2y^2}{2}-\frac{x^3}{3}+\frac{18y^2-6y+1}{108}\right)=C"